OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number.

OpenStudy (mertsj):

\[y=kxn^2\]

OpenStudy (anonymous):

its 260

OpenStudy (anonymous):

y / (xn^2) = k 25 / (3•16) = 25/48 = k y / (5•100) = 25/48 y = 500 • 25/48 y = about 260

OpenStudy (mertsj):

So what is your question then?

OpenStudy (anonymous):

I didnt know it at first then I tried it